If an average family of 4 goes to fast-food restaurants 3 times a week (let’s say Wednesday, Friday, and Saturday for dinner), and they spend $25 each time, how much will they spend over the course of the month? 



1
Expert's answer
2021-03-17T07:01:12-0400

If they visit the restaurant 3 times a week, and 4 weeks makes a month z then


3 times = 1 week

x times = 4 weeks


=> x = 4 × 3 times = 12 times


Since they spend $25 each times.

Total money spent in a month = 12 × 25 = $300